Circular No. 7049

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA OPHIUCHI 1998 R. J. Rudy, T. K. Tessensohn, S. Mazuk, C. J. Rice, and D. K. Lynch,

Aerospace Corporation; and R. C. Puetter, University of California at San Diego, report 0.8–2.5-μm spectroscopy of N Oph 1998 using the Aerospace near-infrared imaging spectrograph and the 3-m Shane telescope at Lick Observatory on June 19.40, June 21.37, and Oct. 2.15 UT: “On the first night, the Paschen and Brackett emission lines dominated the spectrum, although He I line at 1.083 μm was the strongest single line present, and He II was very weak. There was also an unusual continuum emission feature present between 1.07 and 1.10 μm that was about twice the height of the surrounding continuum. Three nights later the nova’s rapid evolution to higher excitation conditions was evident from the much stronger He II lines that had increased by a factor of four relative to the hydrogen lines. The C III line at 0.971 μm also appeared the second night, a line that we have heretofore seen only in Wolf-Rayet stars and that almost certainly indicates an overabundance of carbon. The strange continuum feature was also present. About 100 days later on Oct. 2.15, the emission lines were very weak (5–15 percent of the continuum), and only a few Paschen features and lines of He II were present. The unusual continuum feature was absent. In both June and October the continua, though somewhat different in shape, were blue and resembled a stellar continuum. There was no evidence of a long-wavelength upturn indicative of thermal emission from dust. On the three dates in June and October, the continuum magnitudes at J (1.25 μm) were 9.4, 10.9, and 15.0, respectively.”

V4334 SAGITTARII W. Liller, Instituto Isaac Newton, Viña del Mar, Chile; H. W. Duer-

beck, Space Telescope Science Institute, Baltimore; and A. van der Meer and A. van Genderen, Leiden University, report that this object (cf. IAUC 6825) has undergone another dramatic fading due to dust formation. V magnitudes by Liller (L; 0.2-m telescope) and by Duerbeck and van der Meer (D and M; ESO Dutch 0.9-m telescope): Aug. 12.02 UT, 12.06 (L); 13.99, 12.47 (L); 18.02, 12.56 (L); Sept. 12.98, 13.93 (D); 15.00, 13.91 (L); 17.98, 14.13 (L); Nov. 8.01, 15.11 (M); 9.03, 15.15 (M). Average colors for the last two nights are U − B = +1.95, B − V = +2.52, V − R = +1.68, V − i (Gunn i) = +3.15. On eight nights between Sept. 19.0 and Oct. 29.0, the object was fainter than V = 15.0 (L).

1998 November 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green